find the slope-intercept equation that passes through (-1,2) and is perpendicular to the line with the equation 6y-x=12
Found 2 solutions by mananth, Solver92311:
Answer by mananth(16946)


6y-x =12
6y=x+12
y= (1/6)x +12
slope = 1/2
Slope of perpendicular line will be negative reciprocal of (1/6)
slope = -6
This line with slope -6 passes through -1 ,2
plug these values to get value of b
y=mx+b
2= -6*-1 +b
2+6=b
b=-4
The required equation is y = -6x -4

Answer by Solver92311(821)



Solve the given equation for in terms of . The coefficient on will then be the slope of the graph of that equation.

Perpendicular lines have negative reciprocal slopes. Find the negative reciprocal of the slope of the given line which will be the slope of the desired line.

Once you have the slope of the desired line, use the Point-Slope form of an equation of a line to derive the desired equation.

Point-Slope Form



Where is the given point and is the calculated slope.


John

My calculator said it, I believe it, that settles it

From
I > Ø